### v1

# Promotion of Scientific Publications on ArXiv and X Is on the Rise and Impacts Citations

Abstract
In the evolving landscape of scientific publishing, it is important to understand the drivers of high-impact research, to equip scientists with actionable strategies to enhance the reach of their work, and to understand trends in the use modern scientific publishing tools to inform their further development. Here, based on a large dataset of computer science publications, we study trends in the use of early preprint publications and revisions on ArXiv and the use of X (formerly Twitter) for promotion of such papers in the last 10 years. We find that early submission to ArXiv and promotion on X have soared in recent years. Estimating the effect that the use of each of these modern affordances has on the number of citations of scientific publications, we find that in the first 5 years from an initial publication peer-reviewed conference papers submitted early to ArXiv gain on average $21.1\pm 17.4$ more citations, revised on ArXiv gain $18.4\pm 17.6$ more citations, and promoted on X gain $44.4\pm 8$ more citations. Our results show that promoting one’s work on ArXiv or X has a large impact on the number of citations, as well as the number of influential citations computed by Semantic Scholar, and thereby on the career of researchers. We discuss the far-reaching implications of these findings for future scientific publishing systems and measures of scientific impact.
Introduction
Scientific publications started in the $17^{th}$ century when decisions about whether a paper would be published were made by a select group of experts. As scientific publications proliferated in the $20^{th}$ century, editors of these journals alone could no longer make informed decisions about publications for such a large volume of papers. This led to peer review becoming a standard practice in academia. In the 1990s, the advent of the internet greatly helped with the dissemination of scientific publications. Modern science offers affordances such as publishing preprints before peer-review, revising preprints, promoting scientific works on social media, and viewing researcher profiles listing all co-authored publications and the overall number of citations. However, it is not clear what are the trends in scientific publishing and evaluation, what is their impact, and what will the future bring.
Open-access e-print repositories such as ArXiv and SSRN, and publishing platforms such as OpenReview.net and PeerJ are at the forefront of this change (Sutton and Gong 2017; Abdill and Blekhman 2019), facilitating swift exchange of scientific information. ArXiv hosts 2 million e-prints in various technical fields, while SSRN houses 1.3 million papers for social sciences. These platforms enable early publications and later revisions of unreviewed manuscripts. Then, once a preprint is published, the work can be instantly promoted on social media. Early preprint publication and promotion on social media likely contribute to the number of citations of respective papers, which arguably is the most popular measure of scientific impact. This phenomenon incentivizes scientists to publish early preprints and promote them on social media, without any peer review of their works, suggesting that the publication of preprints and promotion on social media will become standard practices, if not the case already.
Unfortunately, these modern affordances of scientific publishing process, can lead to the spread of questionable scientific results. For instance, in July 2023, a Korean team claimed in two ArXiv publications the discovery of a room-temperature superconductor, which quickly became widely discussed on social media. By mid-August a series of papers from major labs debunked the claim (Johnson 2023). Prior works suggest that social influence make it more challenging to distinguish between high and low quality content (Salganik, Dodds, and Watts 2006), including in the context of scientific publications (Fisher and Parisis 2015; Baddeley 2015; Resnik and Smith 2020), and recent research introducing novel measures of scientific impact discusses a correction for such social impact (Ke, Gates, and Barabási 2023).
In this context, it is not surprising that not all venues embrace early unreviewed scientific manuscript publications. Some major computer science conferences, enforce anonymity periods on ArXiv submissions before conference submission to uphold double-blind review policies. Their critics argue this might decrease paper visibility and subsequent citations.
Figure 1: Top-30 computer science conferences by paper count in our cross-referenced dataset. The colors correspond to research areas in computer science according to Sutton and Gong (2017).
Despite these potential implications of early preprint publications and their promotion on social media, there is a lack of comprehensive, large-scale studies on the use of these affordances and their impact on the number of citations. This study bridges this gap and reflects on the future of scientific publishing. The study focuses on the field of computer science due to its rapid adoption of new scientific publishing technologies. In the field of computer science, ArXiv and X are currently the go-to platforms for publishing preprints and disseminating research findings. Specifically, we address the following research questions.
RQ1: Are early submissions and revisions on ArXiv, and the practice of tweeting about research publications, emerging as the de facto standards?
ArXiv offers next-day publication that can lead to a first-mover advantage and boasts of over 5 million monthly active users (Tech 2023). However, it is unclear whether publishing first on ArXiv, rather in a conference proceeding, impacts the number of citations.
RQ2: What is the effect of publishing a paper on ArXiv before publishing it in conference proceedings on the number of citations of such a paper?
ArXiv also offers the possibility of revising one’s work. On the one hand, a revision may lead to a better publication, especially after receiving expert feedback. On the other hand, if the paper was not fully developed at the time of preprint publication, this could negatively affect its perception among scientists and the number of citations it will receive.
RQ3: What is the effect of revising a paper on ArXiv on the number of citations of such a paper?
Next, we investigate the effect of dissemination of scientific research on X.
RQ4.1: Can promoting a paper on X (formerly known as Twitter) result in more citations of the paper?
RQ4.2: Is the effect of tweets by paper authors different than that of tweets by other users?
Finally, we hypothesize that more modern measures of scientific impact may be less susceptible to influence of social media, following the intuition that important citations correspond to a deep semantic connection between the citing and cited papers.
RQ5: Do newer methods for judging article-level research impact, particularly Semantic Scholar’s highly influential citations, show reduced sensitivity to social exposures on X in comparison to the traditional number of citations?
These research questions have not been sufficiently addressed yet, since it is challenging to answer them quantitatively. Performing a controlled experiment across multiple conferences over multiple years through intervention is intractable. Calculating treatment effects using observational data is difficult because it requires bridging mutliple datasets. To this end, we need information about peer-reviewed conference publications, ArXiv publications, relevant posts on X, and fine-grained information about the dates of publications. Another challenge is that there are multiple factors that could confound treatment effect estimates.
In this paper, we study publications in computer science conference proceedings across multiple areas. Computer science has a unique publication system, prioritizing conferences over journals. Precise publication dates are crucial for measuring trends in early preprint submissions to ArXiv (RQ1) and for estimating its effect (RQ2). While we utilized WikiCFP to gather conference dates, the same level of precision was not attainable for journal publications, which is the prominent mode of publication outside of computer science. Historically, the field of computer science has been relatively quick to embrace emerging technologies for scientific publication, facilitating the creation of datasets necessary for conducting this study, such as WikiCFP. To answer the research questions addressed in this paper, we create a dataset by cross-referencing five data sources: Semantic Scholar, ArXiv, WikiCFP, Altmetric.com, and X, which includes the largest and most prestigious computer science venues (Fig. 1). The dataset contains 18113 papers published over 34 years.
To account for confounding biases in the dataset, we rely on treatment effect estimation techniques, including state-of-the-art doubly robust estimators. We define treatments that address our four main research questions (RQ2-5) and study their effects on the number of citations of a paper (RQ2-4), or influential citations (RQ5), in the 5 years since the date of first publication (on ArXiv or conference proceedings). We use the terms ArXiv-first effect (RQ2), revision effect (RQ3), effect of tweeting (RQ4.1), and effect of author(s) tweeting (RQ4.2) to refer to these effects in the remainder of this paper. We estimate the effects while controlling for potential confounders, such as longer time to accrue citations and self-selection of better papers in hot research areas to be posted on ArXiv before conference publication.
Our study has implications for future publishing models and measures of scientific impact, as we discuss in the last section of this manuscript, which are particularly relevant given that the computer science community is uniquely positioned to lead the development of novel publishing models and scientific impact measures.
Related Work
While extensive literature has studied characteristics of open access publications, described next, in this study we investigate a distinct set of questions. Our inquiry centers on the impact of submitting earlier to ArXiv than to a peer-reviewed venue, and subsequently revising. Additionally, we investigate the effect of promoting research on social media. While Feldman, Lo, and Ammar (2018) have explored the ArXiv-first effect (RQ2) in a smaller study without doubly robust estimation, the effects outlined in RQ1 and RQ3-5 remain unexplored or underexplored.
Trends in scientific publishing (RQ1)
Publishing on ArXiv and promoting a paper on X have increased in popularity in the computer science community. Sutton and Gong (2017) observed a remarkable surge in ArXiv publishing within computer science, reporting that 23% of all papers in 2017 were published on ArXiv (among which 56% were preprints), compared to just 1% in 2007. We study trends in early ArXiv submissions, revisions, and social media promotion for computer science publications from 2013 to 2022.
Publishing preprints on ArXiv (RQ2) and the open access effect
We measure the effects of submitting a version of one’s paper to ArXiv before the conference, i.e., the ArXiv-first effect. Elazar et al. (2023) also studied early ArXiving but focused on its effect on conference acceptance rather than citations which is our focus. Feldman, Lo, and Ammar (2018), on the other hand, performed an analysis on citations and found that there is a 65% increase in citations that can be attributed to the ArXiv-first effect. This study, however, analyzed only the top 16 conferences, used a subset of the confounders we take into account, and did not apply doubly robust estimation.
Multiple prior works find that open access (OA) papers have higher number of citations (Harnad and Brody 2004; Antelman 2004; Cheng and Ren 2008; Fu and Hughey 2019; Alkhawtani, Kwee, and Kwee 2020). This relation was thought to be causal until a study by Kurtz et al. (2005) about research articles in astrophysics on ArXiv showed that this difference in the number of citations can be explained by two sources of bias that were previously unaccounted for. First, the early view bias refers to the situation where articles that are submitted early to a preprint repository, before peer-review publications, have more time to accrue citations. Second, the self-selection bias refers to the phenomenon that better papers are selected to be published in preprint repositories in additoin to peer-reviewed venues. Moed (2007) states that the self-selection bias can occur due to two reasons: i) the author is very well-reputed and is confident about putting out unfinished work (established author bias), ii) the author, irrespective of their reputation in the field, publish their better works on ArXiv (quality bias). These papers in some cases can be unfinished and the authors keep revising them with time. Controlling for these two biases makes the OA advantage negligible for papers in astrophysics. Since then, there have been multiple studies in different disciplines that study these effects. Davis and Fromerth (2007) find that for mathematics papers there is no significant early view effect and the difference in citations between OA and non-OA can be explained by the self-selection bias. More specifically, they find that there is a quality differential between papers that get submitted to ArXiv compared to papers that do not.
While we do not directly compute the OA effect (comparing the impact of submitting to ArXiv against not doing so) in this paper, we leverage insights from these studies to inform our work. We control for the early view bias by only considering citations over a fixed 5-year window after the date of first publication, independently whether it is on ArXiv or in conference proceedings. We also control for self-selection bias using a combination of author characteristics and conference characteristics to lessen the influence of established author and quality biases.
Revising preprints (RQ3)
We explore the impact of paper revisions on citations (termed the ”revision effect”). The prior research on this topic is scarce. Rigby, Cox, and Julian (2018) examined the correlation between revisions and the number of citations for a single social sciences journal. We perform a study at a much larger scale for computer science publications and use treatment estimation techniques to account for confounders.
Dissemination on social media (RQ4)
The internet has profoundly changed how scientists share their research, with platforms like X now serving as essential tools for dissemination of scientific information. According to Wang, Chen, and Glänzel (2020), preprints attract more attention on social media, expediting scholarly communication. Numerous studies reveal positive correlations between mentions on X and paper citations (Eysenbach 2011; Sudah et al. 2022; Klar et al. 2020), underlining the influence of online platforms in academic impact.
In a randomized control trial (RCT) by Luc et al. (2021), 112 cardiothoracic surgery articles from leading journals were randomized 1:1, resulting in an average of $3.1\pm 2.4$ citations for tweeted articles after 1 year, compared to $0.7\pm 2.4$ for non-tweeted ones. Conversely, Tonia et al. (2016) conducted a similar RCT for the International Journal of Public Health (IJPH) and found no differences in citations or article downloads between the exposure and control groups.
Gunaratne, Haghbayan, and Coomes (2020) go beyond analyzing the impact of tweeting. They investigate whether authors tweeting about their papers in pulmonary and critical care medicine receive more citations than tweets from other accounts. The authors find that tweets from paper authors lead to 1.41 times more citations in one year and 1.51 times more in total. In a similar study on shoulder and elbow surgery publications, Sudah et al. (2022) find no significant increase in citations when papers are tweeted by the authors.
The effect of tweeting remains unexplored in computer science. We quantify the impact of tweeting on the number of citations and examine the impact of authors tweeting about their work versus tweets from other accounts.
Novel measures of scientific impact (RQ5)
We also study the effects of our treatment on more contemporary measures of paper-level impact such as highly influential citations (Valenzuela-Escarcega, Ha, and Etzioni 2015). Semantic Scholar employs a machine learning algorithm to calculate this metric, which assesses a citation’s significance by analyzing contextual information extracted from the paper’s content. The features used to classify a citation as highly influential include the text right before the citation, the section where the citation occurs, author overlap between the citing and the cited paper, etc. Therefore, the total number of highly influential citations may be a better measure of research impact. Thus, we hypothesize that they are relatively less affected by promotion on X and ArXiv than the plain number of citations.
Data
Figure 2: Step-by-step description of our dataset development process. “SS” stands for Semantic Scholar. The colored boxes stand for input/output data and the white boxes depict actions taken to produce the output.
We create our dataset using data from five sources (as depicted in Fig. 2). In this section, we describe the details of our dataset development process. Then, we describe the distribution of the number of citations of papers belonging to that dataset and how we use the dataset to answer our research questions.
Dataset development
We have created our dataset using data from five sources: Semantic Scholar 111www.semanticscholar.org (CC BY-NC license), WikiCFP 222www.wikicfp.com, ArXiv 333arxiv.org (CC0: Public Domain license), Altmetric.com, and the academic X API. Semantic Scholar is a search engine for scientific literature developed by the Allen Institute of AI. We used their publicly available dataset to get information about articles, authors, citations, and highly influential citations. WikiCFP is a website that contains details about calls for papers for scientific and technological conferences. We use the data from WikiCFP as a source of important dates relating to the submission and decision deadlines for a conference. To access information about paper submission and revision dates, we utilize the dataset provided by ArXiv. We examined citations spanning a five-year timeframe from the date of first publication (either on ArXiv or in a conference), incorporating papers up to 2017 in our analysis  [^4] Association for Computational Linguistic’s (ACL) anonymity policy started in October 2017 so we have safely ignored that and have not taken the anonymity policy into account in our analysis., as we used Semantic Scholar data up to 2022. Altmetric.com provided us with research access to their proprietary dataset that serves as the data and metadata source for tweets related to each research paper in our dataset. Utilizing the unique tweet ID provided by Altmetric.com, we retrieve the tweet, its metadata, and the author’s metadata using the X API.
Cross-referencing Semantic Scholar and ArXiv.
The ArXiv metadata of computer science papers is cross- referenced with the data from Semantic Scholar by matching the ArXiv IDs or the titles of the articles.
Cross-referencing WikiCFP using conference names.
We use WikiCFP as a source of the following 5 dates: the submission deadline, the notification date (the date when the author is notified about acceptance or rejection), final version (camera-ready) due date, and conference start and end dates. Some important conferences in computer science (like NeurIPS, AAAI, and ACL conferences) did not have all the dates on WikiCFP. We gathered all conferences from CSRankings.org and found that 8 of the top conferences  [^5] The names of these conferences can be found in the Appendix B section. were missing from WikiCFP which we added to our dataset. In addition to these 8 conferences, we identified around 400 conference editions with missing or incomplete WikiCFP information. We annotated their dates manually, using information publicly available on the internet, to include as many important computer science venues as possible and to increase our dataset size. Despite this, we could not find all 5 required dates for around 150 important conferences, e.g., some of the older NeurIPS editions. Overall, we have 46.5k conference editions with full information about them. This data is included in our dataset release  [^6] Code and dataset release: http://bit.ly/48OSnE1.
We matched conference names between the WikiCFP and Semantic Scholar datasets to add information about conference dates for the articles on Semantic Scholar. We noticed a lot of variability in conference names between the two datasets as there are no naming conventions that are strictly followed in either of the two datasets. We extended the conference name matching algorithm used by Demetrescu et al. (2022), but unlike their approach, we do not rely on DBLP for conference names. Instead, we utilize conference names and acronyms from WikiCFP , and conference names from Semantic Scholar. Appendix C contains the details of our algorithm. Our final dataset after cross-referencing Semantic Scholar, WikiCFP, and ArXiv contains 18113 articles with author and citation information (including highly influential citations), conference dates, and ArXiv submission data and is used for calculating trends for ArXiv-first submissions and revisions (RQ1) and to calculate the ArXiv-first and revision effects (RQ2,3).
Figure 3: Comparison of the probability distribution for citations in 5 years after the first date of publication for all computer science papers on Semantic Scholar (SS), those on ArXiv cross-referenced with Semantic Scholar, and our dataset derived from cross-referencing ArXiv, Semantic Scholar, and WikiCFP.
Identifying papers with tweets.
The Altmetric.com API provides rich information about papers that are publicized on social media, including aggregate information (for e.g., total number of posts about a paper on social media), as well as more fine-grained information about each social media post describing a paper. Using the API, we found 13564 papers that have at least one tweet associated with them, 2929 papers that have no tweets associated with them, and Altmetric.com did not contain information about 1620 papers in our dataset.
Identifying papers tweeted by authors and aggregators.
In this section, we discuss the algorithms used to identify papers tweeted by authors and aggregators. We define aggregators as X feeds curating research papers, including both bot accounts using keyword matching or other specialized algorithms on ArXiv and individuals curating papers in their X feeds. We aim to isolate the impact of tweeting (RQ4.1) itself from the distinction between whether the tweet originated from an author or an aggregator. Hence, when computing the treatment effect for tweeting, we factor in whether the paper was shared by an author or an aggregator. Also, to calculate the effect of author(s) tweeting (RQ4.2) we need information about the papers tweeted by authors. For this purpose, we must identify aggregator accounts and employ a name-matching algorithm to associate names between X users and paper authors.
  • Identifying aggregator accounts. To identify aggregators, we set a threshold based on the frequency of tweeting research papers after a close manual inspection, classifying accounts exceeding one paper every 10 days as aggregators. We integrate data on these aggregators’ tweets and followers as a confounding variable in our analysis.
  • Matching authors names to X profiles. We identify papers that authors tweeted about by cross-referencing author names with X users who tweeted about a paper. We employ a name matching algorithm and estimate false negatives using an independent validation classifier. Our name matching algorithm identifies 1,344 papers tweeted by authors, with a true positive rate of 94%. Employing a method from Mane, Srivastava, and Hwang (2005), we estimated approximately 170 false negatives, constituting about 1% of papers in our dataset, which we conclude is negligible. Appendix D contains details on the algorithm and false negative estimation.
This dataset is used to find trends in tweeted papers over the years (RQ1) and to calculate the effects of tweeting (RQ4.1, 4.2).
Dataset characteristics
The dataset we released contains 18113 articles with author and citation information (including highly influential citations), conference dates, and ArXiv submission data. This cross-referenced dataset has a different citation distribution with a higher number of mean and median citations compared to all computer science papers in Semantic Scholar or ArXiv, since only the higher-tier conferences appear in WikiCFP (compare red and green lines in Fig. 3). Our estimates of treatment effects are for papers that are published in such high-tier conferences. We note that for lower-tier conferences, treatment effects will be proportionally lower. Finally, we notice that the number of influential citations is about 7 times smaller than the number of plain citations (compare dashed green and continuous lines in Fig. 3).
Dataset use for treatment estimation
We use all 18113 papers to study the ArXiv-first effect (RQ2) and revision effect (RQ3), and 16493 (=13564+2929) papers for which we can extract tweet information from Altmetric.com to study the effect of tweeting (RQ4.1). To examine the impact of author(s) tweeting as opposed to receiving tweets from a different source (RQ4.2), we focus on the subset of 13564 papers that have associated tweets. In RQ5, we use the same dataset for confounders and treatment variables as in RQ2-4, with highly influential citations serving as the outcome variable.
Methodology for treatment effect estimation
This section describes the control variables utilized in our study and the rationale behind using them, followed by details about feature selection and the models used for treatment effect estimation.
Estimates of effects rely on model specification. If model specification is correct, then estimates tend to be unbiased. Doubly robust methods estimate both the model of treatment and outcome. If one of the models is correct, that is, it matches reality, then the estimate is unbiased. Here, to increase the probability that our estimates of average treatment effects are unbiased, we use doubly-robust methods such as Targeted Maximum Likelihood Estimation (TMLE) (Schuler and Rose 2017). Furthermore, we use multiple model specifications and train a superlearner (Van der Laan, Polley, and Hubbard 2007) for our treatment and outcome models. In addition, to capture the distribution of treatment effects among the papers, we perform standardization using the superlearner outcome model.
Confounding Variable Reason to Control
ArXiv-first effect
Top author: Total number of citations
Top author: h-index Self-selection bias
First author: Total number of citations (established author)
First author: h-index
Number of authors
Conference where the paper was Self-selection bias
accepted (quality)
Conference rating
Was the paper tweeted? Self-selection bias
Subject (CS, Physics etc.) Area bias
Computer Science subfield (AI,
Networking, etc.)
Publication date Temporal bias
Revision effect
+Paper submitted to ArXiv before Revision Effect
or after conference date might be correlated
with ArXiv-first effect
Effect of tweeting
+Sum of followers of X users who Audience size
tweeted about a paper
+Tweeted by author Exposure through
+Tweeted by aggregator authors or aggregators
on X
+Was the paper revised on ArXiv? Effect of tweeting
might be correlated
with Revision
-Was the paper tweeted? Removing the treatment
variable
Effect of author(s) tweeting
+Sum of followers of authors on X Audience size
+Sum of followers of aggregators on X
+Tweeted by scientist Exposure through
+Tweeted by science communicator other channels
+Tweeted by practitioners on X
+Tweeted by members of the public
-Sum of followers of X users who Audience size already
tweeted about a paper captured using
the above confounders
-Tweeted by author Removing the treatment
variable
-Tweeted by aggregator
Table 1: Confounding variables and the reason to control for the confounding variable. The set of confounders under ArXiv-first effect is common for all treatments. We cumulatively add or remove new confounding variables with a ”+” or ”-” in front of their names, respectively.
Figure 4: When do ArXiv and X or Twitter publications happen? Histogram of the number of days since (a) the conference submission date, (d) the camera-ready deadline to the first ArXiv submission, histogram of the number of days since (b) the conference submission date, (e) the camera-ready deadline to ArXiv revisions, and histogram of the number of days since the (c) the conference submission date, (f) the camera-ready deadline to the first tweet by an author.
Our outcome variables here are the number of total citations (RQ2-4) and the number of highly influential ones (RQ5). We only consider citations for five years since the date of first publication (on ArXiv or in the conference). By doing this, we can compare older papers, that might have higher citation counts as they were published earlier, directly with newer ones, that is we control for the early access. To estimate the treatment effects for our hypotheses, we control for potential confounders. Table 1 details all control variables. First, we describe the core set of control variables, which is common for calculating all the effects (RQ2-4). Then, we describe the adjustment in control variables for calculating the effect of tweeting (RQ4).
Control variables
The self-selection postulate (Kurtz et al. 2005) states that better papers or papers from established authors are published more on ArXiv. This is referred to as the paper quality bias and the established author bias in this paper, respectively. To control for the established author bias, we use the total citation counts and h-index of the first author and the top author (selected based on maximum total citation counts). We use the top 150 conferences (estimated based on the number of papers in each conference in our dataset) as a one-hot encoded feature. We also control for the average number of citations of papers published per conference, and the conference rating from the GII-GRIN-SCIE Conference Rating Initiative  [^7] https://scie.lcc.uma.es:8443/. These confounding variables give us a signal about the quality of the paper and hence can be used to control for the quality bias.
We note that tweets by publication authors can be treated as self-selection signals. Thus, we use the presence of a tweet about a paper as another control variable. We fetch total tweet counts from Altmetric.com for each paper and encode it into a binary variable indicating whether a paper was tweeted or not. Eysenbach (2011) finds that majority of tweets are made on the day of publication or the day after. Therefore, knowing that a paper was tweeted (i.e., if the tweet count is greater than zero) is a good indicator that a paper was tweeted right after publication and encompasses both axioms of the self-selection bias (i.e., papers may be tweeted by an established author and/or due to their quality).
We also use information about confounding variables like the subject category of the paper other than computer science (like physics, biology etc.) if it exists, the computer science subfield it belongs to (like artificial intelligence, databases, etc.), and the conference it was published in to control for biases arising from certain computer science areas being more popular at different periods of time and hence accruing more citations. We call this area bias. The publication date is also used as a control variable to adjust for temporal changes.
Differences in control variables across treatments
When estimating the tweeting effect (RQ4), we control for a few additional variables. We describe these variables and the need to control for them in this section. First,we control for the total number of followers of X users who tweeted about the paper. In this way, we control for the impact of differing X audience sizes. When we calculate the effect of tweeting by authors, we also control for the sum of followers of the authors and aggregators to control for the same bias. Altmetric.com also provides demographic information about people tweeting about a paper dividing X users into scientists, science communicators, practitioners, and members of the public. We use four binary variables to indicate whether the paper was tweeted by each of these demographic groups and use this to control for exposure through other channels except authors on X.
Feature selection
We perform feature selection using Variance Inflation Factor (VIF) to remove multicollinearity among any of the features in the model. We remove the average citations for a conference as a feature as it is collinear with other features.
Models
We use standardization and Targeted Maximum Likelihood Estimation (TMLE) using superlearners for estimating average treatment effects. The treatment variable is always binary and is estimated using a superlearner trained on a mixture of a logistic regression model, Discrete Bayesian Additive Regressive Trees (DBARTS) and Generalized Additive Models (GAM). The outcome model to estimate the number of citations is also a superlearner that utilizes Bayesian Additive Regressive Trees (BART), and random forest regression. We use TMLE (Gruber and Van Der Laan 2009) because it is semiparametric and makes minimum assumptions of the distribution of the data . It is also doubly robust and the targeting step in TMLE optimizes the bias-variance tradeoff for the average treatment effect instead of the bias-variance tradeoff for the outcome which is common in other methodologies. We use standardization on the outcome model to estimate the treatment effects for individual papers and understand the heterogeneity of each of the effects across papers.
Figure 5: Percentage of papers that are submitted first and revised on ArXiv, and tweeted on X.
Results
Treatment TMLE Standardization
ATE p-value Cross-val AUC ATE p-value Train Cross-val
$R^{2}$ $R^{2}$ $R^{2}$
All citations
ArXiv-first effect $21.1\pm 17.4$ 0.02 0.14 0.77 $9.0\pm 3.7$ $<10^{-6}$ 0.49 0.19
Revision Effect $18.4\pm 17.6$ 0.04 0.13 0.73 $12.0\pm 3.6$ $<10^{-6}$ 0.48 0.14
Effect of Tweeting $44.4\pm 8.0$ $<10^{-6}$ 0.19 0.97 $11.9\pm 1.9$ $<10^{-6}$ 0.62 0.26
Effect of Author(s) $28.4\pm 11.2$ $<10^{-6}$ 0.18 0.99 $14.8\pm 8.9$ 0.001 0.61 0.24
Tweeting
Highly influential citations
ArXiv-first effect $3.4\pm 2.9$ 0.02 0.07 0.76 $1.4\pm 0.8$ $2\times 10^{-4}$ 0.43 0.10
Revision Effect $2.5\pm 4.0$ 0.21 0.07 0.72 $2.1\pm 0.7$ $<10^{-6}$ 0.44 0.10
Effect of Tweeting $7.6\pm 1.61$ $<10^{-6}$ 0.10 0.98 $2.0\pm 0.4$ $<10^{-6}$ 0.56 0.14
Effect of Author(s) $6.2\pm 2.5$ $<10^{-6}$ 0.11 0.99 $2.6\pm 2.3$ 0.02 0.58 0.12
Tweeting
Table 2: Average treatment effect (ATE) of each of our treatments on all citations and highly influential citations over 5 years after first publication.
Temporal trends in paper promotion (RQ1)
In this section, we discuss the statistical characteristics and temporal trends of papers published in computer science conferences to answer RQ1 based on our rich dataset.
We are interested in understanding how the fractions of papers submitted early or revised on ArXiv change over time. The percentage of papers submitted before the conference to ArXiv has increased from 48.6% in 2013 to 90.1% in 2022 (Fig. 5). In fact, submitting early to ArXiv has become the norm in computer science given the fast-moving nature of the field. The percentage of papers that are tweeted also show a similar trend with an exponential increase in the past few years, reaching 78% in 2022. Conversely, the percentage of papers revised on ArXiv does not exhibit this trend and remains relatively constant with minor fluctuations throughout the years.
We also analyze when researchers make their first submissions to ArXiv and when they submit a revised manuscript on ArXiv. Fig. 4 shows that there are two major peaks corresponding to the periods when researchers make their first submission to ArXiv: one at the conference submission date and the other at the final version due date (camera-ready deadline) (Fig. 4(a), 4(d)). We see a similar trend for revisions, with researchers submitting revised versions of their transcripts much more often around the time of the final version due date than the conference submission deadline (Fig. 4(b), 4(e)), likely after incorporating changes that the reviewers point out in the peer-review process. We also see similar trends for the first tweet by an author(Fig. 4(c), Fig. 4(f)) where there is a peak at the conference submission date and the final version due date but this occurs only because 99.02% of papers tweeted by authors in our dataset receive their initial tweet from an author on the same day as their initial ArXiv submission.
Treatment effects (RQ2-RQ4)
We find statistically significant positive effects for the four treatments using both TMLE and standardization, suggesting that submitting early, revising, tweeting, and the author(s) tweeting lead to a higher number of citations over 5 years after publication. The results are summarized in Table 2.
For the ArXiv-first effect, we notice a positive average treatment effect ($21.1\pm 17.4$ using TMLE and $9.0\pm 3.7$ using standardization), which suggests that submitting to ArXiv before conference deadline on average adds from $9$ to $21$ citations in the first 5 years from the initial publication. It is possible that ArXiv, with its millions of subscribers, provides a visibility exceeding that of conference proceedings. The positive ArXiv-first effect could also indicate a significant first-mover advantage (Newman 2009) where a researcher putting their work out early on ArXiv would get higher number of citations compared to others who wait for the completion of the conference peer-review process, which can take several months, disrupting paper’s timeliness and hotness in a fast-paced research area.
The effect of revisions is also positive ($18.4\pm 17.6$ using TMLE and $12.0\pm 3.6$ using standardization), showing that incorporating feedback from the community makes a paper achieve greater impact. It could also mean that important research papers are revised are improved with time, even post-publication, leading to more citations.
We see a positive treatment effect for tweeting as well ($44.4\pm 8$ using TMLE and $11.9\pm 1.9$ using standardization) which suggests that tweets are an impactful tool for publicizing one’s work in the age of the internet. Research paper showing up on X tend to garner much more citations than papers that are not tweeted.
Among the papers that are tweeted, there is also a positive treatment effect if it is tweeted by an author ($28.4\pm 11.2$ using TMLE and $14.8\pm 8.9$ using standardization), as compared to some other account. This suggests that being active on social media, sharing, and promoting one’s work leads to greater visibility and, ultimately, more citations.
Treatment effects per paper.
Fig. 6 displays the heterogeneous distribution of per-paper treatment effects, calculated through standardization. Here, we could not use doubly robust estimation, since it is designed for average treatment effect estimation. The long-tailed distribution of per-paper treatment effects highlights the diverse nature of these effects, indicating effect sizes vary among papers in our dataset based on individual characteristics. We note that for most papers the effect is positive, although not necessarily significant.
Figure 6: Treatment effects for individual papers using standardization. Black and red corresponsds to all citations and highly influential citaions, respectively. The ATE is marked with a solid line.
Average treatment effects over time.
In Fig. 7 we plot the effect of each treatment in the first year through the fifth year. We see that the ArXiv-first effect and the revision effect increase over the first three years and then plateau in the fourth and fifth years. More interestingly, the effect of tweeting mirrors this pattern indicating that the effect of tweeting is not short-term (i.e., an increase in the number of citations in the first couple of years and then a gradual decrease) but that it actually leads to long-term impact in the five-year time period.
Treatment effects for influential citations (RQ5)
Highly influential citations have positive but smaller effect sizes, given their lower frequency. In our dataset, the average citations are 70.56, with highly influential citations averaging 10.37, resulting in a ratio of 7:1 approximately. This ratio is mirrored in our results (see 2), indicating that promotion on ArXiv and X equally affects highly influential citations. The treatment effects for highly influential citations exhibit also nearly the same temporal trends as raw citation counts (Fig. 6) and have comparable distributions per paper (Fig. 7).
Figure 7: TMLE Average Treatment Effects (ATE) calculated over the first 5 years after publication for (a) ArXiv-first effect, (b) Revision effect, (c) Effect of tweeting, and (d) Effect of author(s) tweeting. Black and red correspond to all citations and highly influential citations, respectively. Error bars correspond to bootstrapped 95% confidence intervals. Results in Appendix E.
Overall, our hypothesis that this novel measure of scientific impact is less affected by tweeting does not hold.
Discussion
In this paper, we look at current trends in computer science publications and the effects of submitting to ArXiv before the conference, revising on ArXiv, and tweeting about a paper.
We find that the practice of publishing on ArXiv before the conference and tweeting about research has increased exponentially over the past few years (Fig. 5), indicating a significant shift in publication dynamics in computer science. We also find positive effects for ArXiv-first submissions, possibly indicating a first-mover advantage (Newman 2009). This presents a challenge, as authors may publish preliminary results on ArXiv to claim ownership, accumulating more citations than more comprehensive studies on the same topic presented at conferences.
ArXiv paper revisions also show positive effects, signaling efforts to address weaknesses or that the topic of study is important enough for the authors to keep working on the paper and improve it with time. These enhancements serve as strong indicators of improved quality, resulting in increased citations.
Our study suggests a positive effect of tweeting, i.e., one’s research gets more attention if it is promoted on the X platform highlighting social effects in dissemination of scientific information. This finding suggests that conferences prohibiting ArXiv submissions diminish the paper’s potential impact on the community. Further research should explore whether this policy disproportionately hinders early-career researchers (as suggested by Saxon (2023)), as they are unable to leverage the advantages of promoting their work on platforms like X, where sharing ArXiv preprints is the norm. Saxon (2023) speculates that the accessibility of social media equalizes opportunities for early-career researchers, reducing the influence of institutional prestige.
Although the affordances outlined in this paper enhance the speed of scientific communication and are gaining popularity, they present challenges to both the scientific review process and the assessment of research impact. Our results suggest that ArXiv is becoming a centralized repository of scientific knowledge with millions of monthly users who are on the lookout for new papers on the platform, which may provide the papers on ArXiv more exposure than conference proceedings. On the contrary, unreviewed low-quality pre-prints on ArXiv can potentially spread misinformation. The current peer review process in computer science does not incorporate any form of review for ArXiv publications, even following conference publications, despite the widespread readership of the ArXiv versions of papers. One potential answer to this challenge is to adopt post-publication peer review (Ford 2013; O’Sullivan, Ma, and Doran 2021), where the peer review process takes place after publication, in contrast to the traditional pre-publication peer review system. This approach not only facilitates the rapid dissemination of scientific information but also fosters meaningful discussions among researchers, thereby encouraging collaborations. If the computer science community adopts a system that allows organic and quicker ways for providing feedback, this could lead to new paradigms in scientific publishing, such as the ones enabled by systems such as OpenReview.net, where reviews can be instant, visible publicly, and responded. We anticipate such systems would quicken feedback loops, while at the same time present challenges for science evaluation.
Moreover, we find that citations and more modern measures such as highly influential citations are volatile and can be easily influenced by promotion on ArXiv and X. Therefore, there is a need for measures of research impact that are impervious to these effects. Proposed citation count normalization methods, like field-normalized or network-normalized counts (Waltman 2016; Ke, Gates, and Barabási 2023; Hutchins et al. 2016), offer alternative metrics that are independent of publication year or field. In contrast, altmetrics, such as the ones reported by Almetric.com or PlumX, directly depend on a paper’s social media attention. Our comprehensive analysis reveals the necessity for developing research impact metrics that incorporate the social aspect of sharing scientific information. This concept is also put forth by Ke, Gates, and Barabási (2023), advocating the exploration of network-normalized metrics on a modified citation graph that accounts for social factors in the dissemination of scientific information. Classifying citations as topical (based on topical similarities between the papers) or non-topical (based on social or professional connections between the authors) in the citation graph drawing from research on social ties (Aiello, Schifanella, and State 2014; Grabowicz, Ganguly, and Gummadi 2016; Grabowicz et al. 2013), offers a promising avenue since topical citations are expected to carry more relevance than non-topical ones when analyzing the research impact of a paper. Employing such a socially-informed measure would further enhance the accurate assessment of scholarly influence.
Nevertheless, we believe any new metric for assessing scientific impact will inevitably fall prey to Goodhart’s law, which states that, ”When a measure becomes a target, it ceases to be a good measure.” In other words, we expect that measures of scientific impact will keep evolving, and be gamed. We hope that this evolution will be driven by a scientific process motivated by studies like the one we developed.
Limitations.
In this paper, we have made a conscientious effort to address numerous confounding variables to minimize biases but there still might exist unaccounted confounding variables which may impact our results. Although we use doubly robust estimators, model specification issues could still introduce bias. The higher mean and median citation counts within WikiCFP, where top-tier conferences are prevalent, might lead to higher effect sizes.
Acknowledgements.
We thank Altmetric.com for granting us access to their dataset and Semantic Scholar for making their dataset publicly available.
APPENDIX
A. Distribution of all citations for each treatment
Figure 8: Box plots for the number of citations for each treatment. The red line in the middle of the box represents the median and the orange triangle represents the mean for (a) Number of citations for all papers, (b) ArXiv-first, (c) Revision, (d) Tweeted, and (e) Tweeted by author.
B. Missing top conferences from WikiCFP
List of top conferences in computer science that are missing from WikiCFP but are listed on CSRankings.org:
  1. 1.
    IMWUT: Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies
  2. 2.
    RTAS: IEEE Real-Time and Embedded Technology and Applications Symposium
  3. 3.
    VIS: IEEE Visualization
  4. 4.
    S&P: IEEE Symposium on Security and Privacy
  5. 5.
    Eurosys
  6. 6.
    USENIX ATC: USENIX Annual Technical Conference
  7. 7.
    SIGSOFT FSE: SIGSOFT Foundations of Software Engineering
  8. 8.
    USENIX Security
C. Algorithm to match conference names between WikiCFP and Semantic Scholar
Algorithm 1 Text normalization
Input: String $s$
Output: Normalized string $s\prime$
SubstringList $\leftarrow$ [’part of’, ’held with’, ’conjunction with’, ’colocated with’, ’collocated with’]
$s\prime\leftarrow$ Convert $s$ to uppercase
$s\prime\leftarrow$ Remove diacritical marks from $s\prime$
$s\prime\leftarrow$ Remove special characters from $s\prime$
$s\prime\leftarrow$ Remove stopwords from $s\prime$
$s\prime\leftarrow$ Remove double spaces from $s\prime$
if any substring $sub$ from SubstringList present in $s\prime$ then
     $s\prime\leftarrow$ Drop trailing tokens in $s\prime$ starting from $sub$
end if
return $s\prime$
Algorithm 2 Loose string match
Input: String $s1$, String $s2$
Output: Percentage match $p$
StringMatchCount $\leftarrow$ 0
s1Tokens $\leftarrow$ Get every word in $s1$ by splitting at every blankspace
s2Tokens $\leftarrow$ Get every word in $s2$ by splitting at every blankspace
for every $token$ in $s1Tokens$ do
     if $token$ in $s2Tokens$ then
         remove $token$ from $s2Tokens$
         StringMatchCount += 1
     end if
end for
p $\leftarrow$ StringMatchCount/len(s1Tokens)
return p
Algorithm 3 Conference name matching
Input: Semantic Scholar venue name $ssNames$, WikiCFP venue acronyms $wikiAcronyms$, dictionary containing WikiCFP acronyms to venue names $wikiAcronymsToNames$
Output: List of matches $match$
match $\leftarrow$ []
for every $ssName$ in $ssNames$ do
     for every $wikiAcronym$ in $wikiAcronyms$ do
         if $ssName$ contains $wikiAcronym$ then
              if len($ssName$) $>$then
                  normalizedSSName $\leftarrow$ textNormalization (ssName)
                  normalizedWikiName $\leftarrow$ textNormalization ($wikiAcronymsToNames[wikiAcronym]$)
                  stringMatchPercentage $\leftarrow$ looseStringMatch (normalizedSSName, normalizedWikiName)
                  if stringMatchPercentage $\geq$ 0.75 then
                       match.append($ssName$, $wikiAcronym$)
                  end if
              else
                  match.append($ssName$, $wikiAcronym$)
              end if
         end if
     end for
end for
return match
To match conference names between between Semantic Scholar and WikiCFP, we first create a dictionary that maps conference acronyms to conference names using data from WikiCFP. Next, we check if there is match between the WikiCFP conference acronym and the Semantic Scholar conference name (typically, Semantic Scholar conference titles incorporate the conference acronym within their names.). If there is a match and the Semantic Scholar conference name has less than 2 words in it (i.e., it contains the conference name and the year, for e.g., ACL 2014), we add the conference name pair to $match$, our final list of matched conference names. If the Semantic Scholar conference name has more than 2 words in it, we perform a loose string match on the normalized strings for the Semantic Scholar and WikiCFP conference names. The algorithms for text normalization and loose string matching are defined in Algorithms 1 and 2 respectively. If the percentage match is greater than 0.75, we add the conference name pair to $match$.
D. Name matching algorithm to find tweets from authors
Name matching algorithm:
We obtain tweet IDs for each paper in our dataset from Altmetric.com, then retrieve tweets and metadata using the X API. We perform fuzzy name matching by first converting Unicode characters to ASCII and standardizing to lowercase, and excluding names shorter than 2 characters (to reduce the number of false positives). Utilizing the fuzzywuzzy Python library, we apply a 95% threshold for matching authors’ last names with X usernames (e.g., John Doe). If no match is found, we repeat the process with screen names (e.g., @johndoe). This yields 1,344 papers with matching author names or X profiles. From this set, a random sample of 50 papers contains 94% true positives for the name matching algorithm.
Calculating false negatives for the name matching algorithm:
To find the number of false negatives with this method, we use the methodology outlined in Mane, Srivastava, and Hwang (2005) where two independent classifiers are used to estimate the number of false negatives for each classifier. The number of false negatives ($\hat{n}_{00}$), i.e., the number of true positives misclassified by both algorithms, can be estimated as
$\hat{n}_{00}=\frac{n_{01}\times n_{10}}{n_{11}},$ (1)
where $n_{10}$ is the number of true positives that are identified correctly by classifier 1 but misclassified by classifier 2, $n_{01}$ is the number of true positives that are correctly classified using classifier 2 but misclassified by classifier 1, and $n_{11}$ is the number of true positives correctly classified by both the classifiers.
To perform this analysis, the second classifier (referred to as the validation algorithm) finds the tweets that start with the word ”our” in our dataset. This follows the intuition that tweets from authors discussing their papers often begin with this word. We also tried other algorithms such as finding tweets that contain the word ”our” instead of strictly starting with it, and tweets that check for other strings like ”we” and ”our paper” but the percentage of true positives using these methods were low. We find that the number of papers for which both the name matching algorithm and the validation algorithm give us a positive result is 213. On manual inspection of 50 of these papers, we find that the intersection has 94% true positives. Therefore, the number of true positives at the intersection of the two algorithms is $213*0.94=200.22(n_{11})$. The validation algorithm has 34 more positives that are not part of the positives for the name matching algorithm. Out of these, 26 are true positives ($n_{01}$). And we already know that the number of positives for the name matching algorithm is 1344 with 94% true positives which makes the number of true positives for the algorithm to be $1344*0.94=1263.36$ ($n_{01}$). We use Eq.1 to calculate the total number of false negatives and $\hat{n}_{00}$ in this case is 154.21 ($\approx$ 154).
Using this analysis, we find that we have a total of 1344 tweets by authors out of which 94% are true positives and we are missing around 154 other true positives that are misclassified by both the algorithms and 26 true positives that are correctly classified by the validation algorithm but misclassified by the name matching algorithm, summing up to a total of 170 estimated false negatives for the name matching algorithm. This number constitutes about 1% of all papers for which we did not find an X profile match, so we conclude that it is negligible.
Treatment Year TMLE Standardization
ATE p-value ATE p-value
First $\mathbf{0.5\pm 0.5}$ $\mathbf{0.039}$ $\mathbf{0.3\pm 0.1}$ $\mathbf{<10^{-6}}$
Arxiv-first Second $\mathbf{2.7\pm 1.7}$ $\mathbf{0.001}$ $\mathbf{1.3\pm 0.4}$ $\mathbf{<10^{-6}}$
effect Third $\mathbf{4.7\pm 3.6}$ $\mathbf{0.010}$ $\mathbf{2.2\pm 0.8}$ $\mathbf{<10^{-6}}$
Fourth $6\pm 6.1$ $0.051$ $\mathbf{2.7\pm 1.4}$ $\mathbf{5.1\times 10^{-5}}$
Fifth $\mathbf{7.2\pm 6}$ $\mathbf{0.017}$ $\mathbf{3\pm 1.4}$ $\mathbf{2.7\times 10^{-5}}$
First $\mathbf{0.5\pm 0.4}$ $\bf{0.040}$ $\mathbf{0.3\pm 0.1}$ $\mathbf{<10^{-6}}$
Revision Second $\mathbf{2.5\pm 1.4}$ $\mathbf{0.008}$ $\mathbf{1.5\pm 0.4}$ $\mathbf{<10^{-6}}$
effect Third 4.5 $\pm$ 3.3 $\mathbf{0.008}$ $\mathbf{2.9\pm 0.8}$ $\mathbf{<10^{-6}}$
Fourth $5.6\pm 5.7$ $0.056$ $\mathbf{3.9\pm 1}$ $\mathbf{<10^{-6}}$
Fifth $4.9\pm 7.4$ $0.194$ $\mathbf{3.8\pm 1.3}$ $\mathbf{<10^{-6}}$
First $\mathbf{1.5\pm 0.2}$ $\mathbf{<10^{-6}}$ $\mathbf{0.6\pm 0.1}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{4.9\pm 0.7}$ $\mathbf{<10^{-6}}$ $\mathbf{1.6\pm 0.3}$ $\mathbf{<10^{-6}}$
tweeting Third $\mathbf{9.9\pm 1.6}$ $\mathbf{<10^{-6}}$ $\mathbf{2.9\pm 0.4}$ $\mathbf{<10^{-6}}$
Fourth $\mathbf{13.9\pm 2.5}$ $\mathbf{<10^{-6}}$ $\mathbf{3.5\pm 0.6}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{14.0\pm 3.2}$ $\mathbf{<10^{-6}}$ $\mathbf{3.0\pm 0.7}$ $\mathbf{<10^{-6}}$
First $\mathbf{1.1\pm 0.3}$ $\mathbf{<10^{-6}}$ $\mathbf{0.6\pm 0.2}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{1.1\pm 1}$ $\mathbf{0.037}$ $\mathbf{1.7\pm 0.6}$ $\mathbf{<10^{-6}}$
author(s) Third $1.9\pm 2.03$ $0.072$ $\mathbf{3.1\pm 1.8}$ $\mathbf{5\times 10^{-4}}$
tweeting Fourth $\mathbf{4.3\pm 3.2}$ $\mathbf{0.008}$ $\mathbf{4.1\pm 3.1}$ $\mathbf{0.009}$
Fifth $\mathbf{13.5\pm 5.5}$ $\mathbf{<10^{-6}}$ $\mathbf{5.3\pm 4.33}$ $\mathbf{0.015}$
Table 3: Average treatment effects for all citations using each treatment variable for the first through fifth year of publication. Statistically significant results ($p<0.05$) are in bold.
E. Average treatment effects for the first through fifth year of publication
Table 3 and Table 4 provide details about the treatment effects over the years for all citations and highly influential citations, respectively.
Treatment Year TMLE Standardization
ATE p-value ATE p-value
First $0.09\pm 0.10$ $0.108$ $\mathbf{0.05\pm 0.03}$ $\mathbf{7.5\times 10^{-5}}$
Arxiv-first Second $\mathbf{0.49\pm 0.36}$ $\mathbf{0.008}$ $\mathbf{0.21\pm 0.08}$ $\mathbf{<10^{-6}}$
effect Third $0.59\pm 0.90$ $0.196$ $\mathbf{0.33\pm 0.21}$ $\mathbf{0.002}$
Fourth $\mathbf{0.98\pm 0.96}$ $\mathbf{0.046}$ $\mathbf{0.38\pm 0.22}$ $\mathbf{0.001}$
Fifth $\mathbf{1.10\pm 0.88}$ $\mathbf{0.015}$ $\mathbf{0.38\pm 0.20}$ $\mathbf{2\times 10^{-4}}$
First $0.04\pm 0.10$ $0.426$ $\mathbf{0.05\pm 0.02}$ $\mathbf{4\times 10^{-6}}$
Revision Second $\mathbf{0.46\pm 0.33}$ $\mathbf{0.005}$ $\mathbf{0.28\pm 0.08}$ $\mathbf{<10^{-6}}$
effect Third 0.73 $\pm$ 78 $0.063$ $\mathbf{0.48\pm 0.15}$ $\mathbf{<10^{-6}}$
Fourth $0.76\pm 1.28$ $0.244$ $\mathbf{0.58\pm 0.20}$ $\mathbf{<10^{-6}}$
Fifth $0.55\pm 1.60$ $0.501$ $\mathbf{0.54\pm 0.27}$ $\mathbf{7.1\times 10^{-5}}$
First $\mathbf{0.32\pm 0.05}$ $\mathbf{<10^{-6}}$ $\mathbf{0.11\pm 0.02}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{1.01\pm 0.16}$ $\mathbf{<10^{-6}}$ $\mathbf{0.30\pm 0.05}$ $\mathbf{<10^{-6}}$
tweeting Third $\mathbf{1.79\pm 0.33}$ $\mathbf{<10^{-6}}$ $\mathbf{0.49\pm 0.09}$ $\mathbf{<10^{-6}}$
Fourth $\mathbf{2.20\pm 0.50}$ $\mathbf{<10^{-6}}$ $\mathbf{0.55\pm 0.12}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{2.11\pm 0.61}$ $\mathbf{<10^{-6}}$ $\mathbf{0.47\pm 0.11}$ $\mathbf{<10^{-6}}$
First $\mathbf{0.18\pm 0.05}$ $\mathbf{3.2\times 10^{-6}}$ $\mathbf{0.13\pm 0.06}$ $\mathbf{2\times<10^{-5}}$
Effect of Second $\mathbf{0.63\pm 0.27}$ $\mathbf{6.8\times 10^{-6}}$ $\mathbf{0.32\pm 0.18}$ $\mathbf{3\times 10^{-4}}$
author(s) Third $0.31\pm 0.47$ $0.187$ $\mathbf{0.61,\pm 0.44}$ $\mathbf{0.007}$
tweeting Fourth $\mathbf{2.43\pm 1.15}$ $\mathbf{3.2\times 10^{-5}}$ ${0.79\pm 0.85}$ 0.0710.071{0.071}0.071
Fifth $\mathbf{1.74\pm 0.80}$ $\mathbf{2\times 10^{-5}}$ ${0.94\pm 0.98}$ 0.0590.059{0.059}0.059
Table 4: Average treatment effects for highly influential citations using each treatment variable for the first through fifth year of publication. Statistically significant results ($p<0.05$) are in bold.